The NetBSD Project

<pk@netbsd.org> January 23, 1996

Overview

• What is NetBSD?

• Goals

• Results

• Future Directions • Organization

What is NetBSD?

• a freely available, portable Unix-like system, derived from the 4.4BSD-Lite distribution from the University of California, Berkeley

• the only free system using the same source tree for many hardware platforms

• distributed with complete source code

• supported by many volunteers

Goals

• provide a complete, stable operating system that both professionals and hobbyists can use

• make more code machine-independent

• support more hardware

• binary compatibility/emulation

• comply more closely with standards (POSIX, RFCs)

• avoid encumbering licenses

Results

• 64-bit support

• emulation architecture

• multicast integration

• MI audio

• generic disk instrumentation

• generalized CCD

• locales

• work started on section 9 man–page section

Results (cntnd)

• new ports this year – vax

– mvme68k

– sun4

• correct prof. Falken’s name in example passwd file.

Current platforms

• alpha (TurboChannel and PCI busses)

• amiga (all machines with a MMU)

• atari (TT and Falcon)

• hp300 (all machines except model 425 and 433 desktops)

• i386/i486/Pentium (PCI, EISA, and ISA busses)

• mac68k (SE/30, II series, some Quadras)

• mvme68k (model 147)

• pc532

Current platforms (cntnd)

• pmax (DECstation 5000 series)

• sparc (Sun 4 and 4c architecture; 4m nearly there; no 4d yet)

• sun3 (Sun 3 architecture; no 3x yet)

• vax (11/750; MicroVAX II; MicroVAX III in progress)

• … and more in progress!

Dead platforms

• da30

Current emulations

• BSD/OS

• FreeBSD

• Interactive

• Linux • SCO

• SunOS

• SVR4

• … and more in progress!

• i386–only – WINE – natively and under Linux emulation.

– VM86

Commercial software reported to work

• AFS client

• DOOM

• FrameMaker

• Informix

• Maple

• Matlab • Mosaic

• Netscape Navigator • Oxford English Dictionary

• SimCity

• WordPerfect

Future Directions

• see ‘Goals’

• symmetric multiprocessing

• journaled file system

• kernel–level support for threads • real–time extensions

Organization

• The NetBSD Foundation

– A framework to:

∗ keep track of development

∗ receive donations to run the project

∗ undertake promotional activities

∗ organizing workshops

• The core group

– system architect

– runs critical services (e.g. CVS tree, mailing lists, GNATS)

– point of contact for contributions • The port masters

– are responsible for their individual ports

– review contributions

– do development

• Mailing lists

– current-users

∗ use the daily source snapshots and provide

feedback on recent changes

– netbsd-bugs (GNATS)

∗ keeps track of unresolved issues, including · bug reports · suggested enhancements

– netbsd-announce

Releases

• NetBSD 0.8, April 1993

• NetBSD 0.9, August 1993 • NetBSD 1.0, October 1994 • NetBSD 1.1, November 1995 • NetBSD-current, daily

• periodic releases planned

Places to look

• To obtain NetBSD, look in – ftp://ftp.netbsd.org/pub/NetBSD

• To get information about, or subscribe to, NetBSD mailing lists, send mail to

– majordomo@netbsd.org

• To ask questions or send comments about NetBSD, send mail to

– netbsd-comments@netbsd.org • To get daily source snapshots

– by FTP, look in ∗ ftp://ftp.netbsd.org/pub/NetBSD/NetBSD-current

– by SUP, look for directions in ∗ ftp://ftp.netbsd.org/pub/sup